The difference of two numbers is 91. THe second is 5 less than 5 times the first. What are the two numbers?

Found 2 solutions by Earlsdon, checkley71:
Answer by Earlsdon(6294)

Let x = the fist number and y = the second number.
1) y-x = 91 "The difference of the two numbers is 91"
2) y = 5x-5 "The second number is 5 less than 5 times the first"
Substitute the y from the second equation into the first and solve for x.
1a) (5x-5)-x = 91 Simplify.
5x-x-5 = 91
4x-5 = 91 Add 5 to both sides.
4x = 96 Divide both sides by 4.
x = 24 and...
y = 5(24)-5
y = 120-5
y = 115
Check:
y-x = 91
115-24 = 91
91 = 91 and...
115 = 5(24)-5
115 = 120-5
115 = 115

Answer by checkley71(8403)

X-Y=91 OR X=91+Y
X=5Y-5
91+Y=5Y-5
Y-5Y=-5-91
-4Y=-96
Y=-96/-4
Y=24 ANSWER.
X=91+24
X=115 ANSWER.
PROOF
115=5*24-5
115=120-5
120=120